The anomaly of f (z) when z is given for an infinite is the same as the anomaly of f (1 / w) at w = 0?
Answer #1

Yes because of as limit of w goes to 0 then 1/w = infinity

hence f(1/w) = f(z) at w=0 and z= infinity